OpenStudy (anonymous):

(PLEASE HELP) A farmer has 1800 feet of fence and wants to enclose a rectangular field with two internal fence parallel to one side, as shown below (drawn) What are the dimensions that would maximize the area? ( I would appreciate just getting started on it and I can do the rest)

OpenStudy (anonymous):



OpenStudy (anonymous):

A-area W*L W=A/L fencing 4W+2L=1800 4A/L +2L=1800 mulptiply by L 4A+2L^2 =1800L dA/dL=o for max/min and solve

OpenStudy (anonymous):

dA/dL=4L-1800=0 for max L=1800/4=450 2L=900 4w=1800-900=900 w=900/4=225 so L=450 W=225

OpenStudy (anonymous):